OpenStudy (anonymous):

Write this decimal in numbers. nine hundred eighty five and sixty five hundredths

OpenStudy (anonymous):

985.65

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

It is 985 + 65 = 985 +0.65 = 985.65 --- 100

OpenStudy (anonymous):

Give me a medal if u wish